Good Evening!
I am trying to solve the following by elimination:
{9/2x+1/2y=17/2
{x-5/2y=9/2
Thanks!
Jim

Answer by stanbon(75887)

solve the following by elimination:
{9/2x+1/2y=17/2
{x-5/2y=9/2
================
Multiply thru each equation by 2 to get rid of the denominators:
================
9x + y = 17
2x -5y = 9
-----------
Multiply thru the 1st by 5:
45x + 5y = 85
2x - 5y = 9
-----------
Add and solve for "x":
47x = 94
x = 2
-----
Substitute into 2x - 5y = 9 and solve for "y":
2*2-5y = 9
-5y = 5
y = -1
----------
Solvution : (2,-1)
=======================
Cheers,
Stan H.
=======================